# A Multi-Objective Evaluation Framework for Analyzing Utility-Fairness Trade-Offs in Machine Learning Systems

### v1

@@ Abstract
The evaluation of fairness models in Machine Learning involves complex challenges, such as defining appropriate metrics, balancing trade-offs between utility and fairness, and there are still gaps in this stage. This work presents a novel multi-objective evaluation framework that enables the analysis of utility-fairness trade-offs in Machine Learning systems. The framework was developed using criteria from Multi-Objective Optimization that collect comprehensive information regarding this complex evaluation task. The assessment of multiple Machine Learning systems is summarized, both quantitatively and qualitatively, in a straightforward manner through a radar chart and a measurement table encompassing various aspects such as convergence, system capacity, and diversity. The framework’s compact representation of performance facilitates the comparative analysis of different Machine Learning strategies for decision-makers, in real-world applications, with single or multiple fairness requirements. The framework is model-agnostic and flexible to be adapted to any kind of Machine Learning systems, that is, black- or white-box, any kind and quantity of evaluation metrics, including multidimensional fairness criteria. The functionality and effectiveness of the proposed framework is shown with different simulations, and an empirical study conducted on a real-world dataset with various Machine Learning systems.
Index Terms:
@@
Multidimensional Evaluation, Multi-Objective Optimization, Pareto Frontier, Utility-Fairness Trade-off
I Introduction
@@
The increasing integration of Machine Learning (ML) systems in day-to-day activities offers significant opportunities, but it also raises critical concerns regarding demographic fairness and equity [1, 2, 3, 4]. Fairness in ML pertains to the ethical imperative of ensuring that algorithms and models do not discriminate, or display bias, against individuals or groups based on lawfully protected attributes such as race, gender, age, and/or other characteristics [5]. Fairness is a multi-faceted and complex concept with nuances directly linked to the situation considered [6, 2]. Consequently, balancing and measuring multiple fairness criteria simultaneously, both at a group and individual level, is a challenging task, resulting in different definitions of fairness appropriate for different contexts [7].
@@
An unintended outcome when optimizing for a balanced treatment between genders is the variations in predictive performance across other groups of protected attributes. Many techniques in real-world scenarios improve fairness at the expense of model utility, with the minimum possible error of any fair classifier bounded by the difference in base rates [8]. This fundamental tension in algorithmic fairness has been previously explored to improve the understanding of model bias and the limits of artificial intelligence [9, 10]. Model biases can be introduced through the optimization of certain objectives, hyper-parameter tuning, or simply due to the inherent characteristics of the datasets used to train the models [11], such as biased training data, sampling bias, label bias, exclusion, or historical bias. For example, if features correlated with sensitive attributes (e.g. gender, race) are considered during training, the models might latch onto these attributes, potentially resulting in demographically biased outcomes.
@@
The selection and modeling of fairness criteria is a relatively new research direction that requires clear definitions on the mathematical expression of demographic equity. Most ML approaches are currently evaluated without considering any fairness criteria [12, 13]. In recent years, some works started using unique levels of fairness and utility, which fail to characterize ML systems at every level of the utility-fairness trade-off, thus limiting subgroup and intersectional evaluation [14]. The situation worsens when multiple fairness and utility criteria come into play, i.e., the deployment of diagnostic tools in a hospital that tends to a heterogeneous community. Management may want to ensure, beyond utility, that multiple fairness criteria are satisfied, such as race, gender, and age aiming to provide fair healthcare for the patients.
@@
On the other hand, the development of ML systems addressing multiple objectives has been extensively studied in the context of Multi-Objective Optimization (MOO) systems. In such cases, performance is evaluated following all possible trade-offs between the individual objectives resulting in an N-dimensional graph. This evaluation strategy can provide a basis to better articulate user preferences in model comparison [15]. Although multi-objective measurements have been used in recent works to incorporate single fairness constraints into developed models, to the best of our knowledge, there are no frameworks that enable a comprehensive comparison of ML systems under multiple utility and fairness criteria [16].
@@
In this work, we present an evaluation framework, supported by MOO principles, for the challenging task of comparing ML systems under multiple utility-fairness trade-offs. This approach allows the comparison of multiple systems in a common multidimensional space, where multiple concurrent fairness criteria, can come into play. Our contributions can be summarized as follows:
@@
  • A model agnostic evaluation framework with a compact yet comprehensive representation, both qualitatively and quantitatively, of multiple utility-fairness trade-offs resulting from the deployment of ML systems, facilitating system performance analysis and comparison.
@@
  • Detailed analysis of the proposed framework and rationale through simulations of typical ML systems on synthetically generated data.
@@
  • An empirical study based on a real-world dataset demonstrating the effectiveness of the proposed framework on a concrete use-case.
@@
The paper is organized as follows: Section II reviews previous works tackling the evaluation of utility-fairness trade-off systems. Section III thoroughly describes the proposed evaluation framework, with a set of use-cases, and MOO principles behind it. The applicability of the framework in a real-world scenario is shown in Section IV, with the analysis of ML systems for a medical imaging task. Finally, a discussion and conclusion with the key points and limitations from this study is presented in Section V and Section VI.
II Background and Related Work
@@
Fairness in machine learning can be categorized based on criteria, sources of bias, perspectives, methodologies, and trade-offs. Fairness criteria include demographic parity, equality of opportunity, equalized odds [17], and predictive parity, each focusing on equitable outcomes or error rates across groups. Sources of bias can stem from data (e.g. underrepresentation), algorithms (e.g. prioritizing accuracy/utility over fairness), or human involvement (e.g. subjective labeling). Perspectives of fairness include individual fairness (similar individuals receive similar predictions), group fairness (equitable treatment across groups), and subgroup fairness (addressing intersectional identities). Methodologies to enforce fairness involve pre-processing data (e.g. balancing representation) [18, 19, 20], in-processing adjustments (e.g. modifying loss functions) [21, 22], and post-processing predictions (e.g. calibration techniques) [17, 23, 24]. Achieving fairness often involves trade-offs, such as balancing it with utility [25, 26], and interpretability [27].
@@
Whereas ML systems are typically developed (and evaluated) using a single utility criteria, they are often deployed in scenarios where multiple objectives must be respected. A modern example of this condition relates to the deployment of ML systems under one or multiple demographic fairness constraints [25, 28, 29]. In this context, we argue evaluation techniques cross-pollinated from multi-objective optimization (MOO) offer a rich set of primitives allowing for a comprehensive performance characterization under multiple criteria that can streamline system evaluation in this realm.
@@
The principal aim of MOO is to ascertain solutions that lie on, or are proximate to, the set of the optimal performance points called the Pareto Front (PF), resulting in a spectrum of ideal trade-offs among the various objectives. This methodology equips decision-makers with the means to select the most favorable compromise amidst conflicting goals, fostering more informed and balanced decision-making [30]. The trade-off selection procedure in MOO is therefore critical and affects the quality of service for the deployed system, especially in the case of conflicting objectives. Assessing the quality of these trade-off systems is comparative and encompasses criteria such as proximity to the Pareto optimal set (convergence), the distribution/spread of the points in the objective space (diversity), and the cardinality of solutions (capacity) [31]. These criteria are evaluated by MOO specific performance indicators that have been studied in previous works (refer Section III-B1 for details) [32, 30, 33, 34].
@@
Even though the modeling of trade-off for demographic fairness-accuracy PF is well-known [9, 35], performance indicators for the quality of the PF have rarely been exploited in the context of fairness. Yang et al. developed a bias mitigation framework that incorporated the Area Under the Curve (AUC) metric, while considering both inter/intra group AUC simultaneously [36]. However, the bias mitigation framework does not provide an evaluation protocol for the utility-fairness trade-off, instead it leverages the AUC to address fairness performance. Little et al. proposed a scalar measure of the area under the curve from the trade-off between fairness and accuracy [16]. The generated curve outlines the empirical Pareto frontier consisting of the highest attained accuracy within a collection of fitted models at every level of fairness. Although, the work in [16] focuses on similar issues as in this study, it does not address the challenge of comparing multiple ML systems in high dimensions. Additionally, the analysis of the PF is superficial, ignoring important performance indicators for diversity and capacity, providing an incomplete evaluation of compared ML strategies. To tackle the aforementioned issues, a more flexible evaluation framework is needed to accommodate different fairness criteria and utility metrics, facilitating a straightforward comparison and analysis of results from different algorithms. The method should be model-agnostic, allowing for real-world comparisons among trade-off systems that may have been optimized using different objectives. Furthermore, since the utility goals of the model across multiple objectives often diverge from fairness goals, performance indicators of the optimal PF solutions can provide a deeper understanding of the trade-offs across these objectives [26]. The proposed evaluation framework bridges the gap between these issues and their solutions, providing a comprehensive guideline on how this can be achieved in the following sections.
III Methodology
@@
In MOO, each individual objective is considered a distinct characteristic that needs to be optimized to its fullest potential. The trajectory of the optimization is determined by objective functions in a cooperative way. Conflicting objectives increase the complexity of optimization forcing cooperation, thus making it harder to achieve optimal solutions. This trade-off is measured at evaluation time by using multiple metrics directly related to each of the objectives. Likewise, utility and fairness can result in conflicting objectives challenging the optimization of both objectives through the same ML strategies. A good level of utility is typically achieved by sacrificing fairness, or through less biased models that might reduce utility.
@@
III-A Use-cases

### v2

@@ Abstract
The evaluation of fairness models in Machine Learning involves complex challenges, such as defining appropriate metrics, balancing trade-offs between utility and fairness, and there are still gaps in this stage. This work presents a novel multi-objective evaluation framework that enables the analysis of utility-fairness trade-offs in Machine Learning systems. The framework was developed using criteria from Multi-Objective Optimization that collect comprehensive information regarding this complex evaluation task. The assessment of multiple Machine Learning systems is summarized, both quantitatively and qualitatively, in a straightforward manner through a radar chart and a measurement table encompassing various aspects such as convergence, system capacity, and diversity. The framework’s compact representation of performance facilitates the comparative analysis of different Machine Learning strategies for decision-makers, in real-world applications, with single or multiple fairness requirements. In particular, this study focuses on the medical imaging domain, where fairness considerations are crucial due to the potential impact of biased diagnostic systems on patient outcomes. The proposed framework enables a systematic evaluation of multiple fairness constraints, helping to identify and mitigate disparities among demographic groups while maintaining diagnostic performance. The framework is model-agnostic and flexible to be adapted to any kind of Machine Learning systems, that is, black- or white-box, any kind and quantity of evaluation metrics, including multidimensional fairness criteria. The functionality and effectiveness of the proposed framework are shown with different simulations, and an empirical study conducted on three real-world medical imaging datasets with various Machine Learning systems. Our evaluation framework is publicly available at https://pypi.org/project/fairical.
keywords:
@@
Machine Learning, Multidimensional Fairness Evaluation, Multi-Objective Optimization, Utility-Fairness Trade-off, Medical Image Analysis
doi:
10.59275/j.melba.2025-ab9a
volume: 3
1 Introduction
@@
The increasing integration of Machine Learning (ML) systems in day-to-day activities offers significant opportunities, but it also raises critical concerns regarding demographic fairness and equity (Xinying Chen and Hooker, 2023; Pessach and Shmueli, 2022; Starke et al., 2022; Rabonato and Berton, 2025). Fairness in ML pertains to the ethical imperative of ensuring that algorithms and models do not discriminate or display bias, against individuals or groups based on lawfully demographic attributes such as race, gender, age, and/or other characteristics (Barocas et al., 2023). Fairness is a multi-faceted and complex concept with nuances directly linked to the situation considered (Castelnovo et al., 2022; Pessach and Shmueli, 2022). Consequently, balancing and measuring multiple fairness criteria simultaneously, both at a group and individual level, is a challenging task, resulting in different definitions of fairness appropriate for different contexts (Dutt et al., 2023).
While Multi-Objective Optimization (MOO) provides a mathematical foundation for balancing objectives in conflict, fairness in ML extends beyond conventional optimization paradigms. Equity encompasses multiple, often conflicting notions, such as group or individual fairness and equality of opportunity, that reflect distinct ethical and social considerations. These criteria cannot be meaningfully represented by a single objective function, as optimizing for one dimension of fairness may amplify disparities in another. In this respect, ML based approaches that model and evaluate the utility-fairness trade-off across multiple objectives can provide valuable insight into how fairness manifests under different operational and demographic conditions. However, as highlighted by Selbst et al. (2019) in five traps, over-abstracting fairness into purely mathematical formalizations can have a risk of detaching it from the reality. Therefore, such analyses should be interpreted as structured, assumption-bound explorations of fairness dynamics, rather than universal solutions to the fairness problem in ML.
@@
An unintended outcome when optimizing for a balanced treatment between genders is the variation in predictive performance across other groups of demographic attributes. Many techniques in real-world scenarios improve fairness at the expense of model utility, with the minimum possible error of any fair classifier bounded by the difference in base rates (Zhao and Gordon, 2022). This fundamental tension in algorithmic fairness has been previously explored to improve the understanding of model bias and the limits of artificial intelligence (AI) (Wei and Niethammer, 2022; Wang et al., 2020). Model biases can be introduced through the optimization of certain objectives, hyperparameter tuning, or simply due to the inherent characteristics of the datasets used to train the models (Yang et al., 2024), such as biased training data, sampling bias, label bias, exclusion, or historical bias. For example, if features correlated with sensitive attributes (e.g., gender, race) are considered during training, the models might latch onto these attributes, potentially resulting in demographically biased outcomes.
@@
The selection and modeling of fairness criteria are relatively new research directions that require clear definitions on the mathematical expression of demographic equity. Most ML approaches are currently evaluated without considering any fairness criteria (Akter et al., 2022; Lu et al., 2020). In recent years, some works started using unique levels of fairness and utility, which fail to characterize ML systems at every level of the utility-fairness trade-off, thus limiting subgroup and intersectional evaluation (Buolamwini and Gebru, 2018). The situation worsens when multiple fairness and utility criteria come into play, as the fairness-aware stakeholders and decision-makers may want to ensure, beyond utility, that multiple fairness criteria are satisfied, such as race, gender, and age aiming to provide fair ML services in production under different demographic settings.
In medical imaging, fairness challenges become more prominent because diagnostic processes often involve multiple and conflicting clinical and operational objectives, as ML models are increasingly used to support decisions in almost every field, such as radiology, ophthalmology, and dermatology. Furthermore, variations in data acquisition conditions, patient demographics (e.g., the deployment of diagnostic tools in a hospital that tends to a heterogeneous community), and disease prevalence can lead to systematic biases in model predictions. For instance, screening models for ophthalmology risk must retain high sensitivity to avoid missed cases, while also ensuring that performance does not systematically degrade for specific demographic subgroups. A representative example arises in glaucoma, an eye disease that exhibits a higher prevalence among Black populations, and within this group, male individuals show greater vulnerability compared to females (Khachatryan et al., 2019; Luo et al., 2024). The high prevalence of glaucoma contrasts with the scarcity of available data from the Black community, contributing to disparities in model performance across racial groups and even between genders within the same group. This mismatch highlights the necessity of evaluating ML models not only for diagnostic utility but also for their equitable behavior across different demographic attributes. Consequently, utility (diagnostic performance) and multiple fairness constraints must be considered simultaneously. A principled way to analyze such conflicting requirements is to employ multi-objective formalizations, where each fairness criterion and the utility metric are treated as separate but jointly optimized objectives rather than in isolation.
@@
On the other hand, the development of ML systems addressing multiple objectives has been extensively studied in the context of MOO systems. In such cases, performance is evaluated by considering all possible trade-offs between the individual objectives, resulting in an N-dimensional graph. This evaluation strategy can provide a basis to better articulate user preferences in model comparison (Gong and Guo, 2023). Although multi-objective measurements have been used in recent works to incorporate single fairness constraints into developed models (Little, 2023), to the best of our knowledge, there are no frameworks that enable a comprehensive comparison of ML systems under multiple utility and fairness criteria. In particular, multiple fairness considerations based on MOO become even more critical in the context of medical imaging. A framework that accounts for multiple fairness constraints simultaneously is therefore essential to ensure equitable diagnostic performance across different patient subgroups characterized by demographic attributes.
Conceptually, fairness in ML can be regarded as a multidimensional evaluation problem rather than a single optimization objective. Instead of focusing on maximizing a particular fairness metric, it is often more informative to examine how ML systems perform across a spectrum of utility-fairness trade-offs. Such an analysis enables a clearer understanding of both ideal cases, where equity across demographic groups is maximized and practical situations in which certain fairness dimensions may be compromised due to data, operational, or design constraints. Considering fairness in this way allows researchers and practitioners to interpret model behavior within an explicit space of trade-offs, rather than through isolated or averaged/maximized performance scores. This perspective establishes the conceptual foundation for evaluating fairness-aware ML systems in a structured, multi-objective manner, before introducing any specific methodological framework.
@@
In this work, we present an evaluation framework, supported by MOO principles, for the challenging task of comparing ML systems under multiple utility-fairness trade-offs. This approach allows the comparison of multiple systems in a common multidimensional space, where multiple concurrent fairness criteria can come into play. The framework’s applicability is demonstrated through use cases based on medical imaging, where fairness constraints play a crucial role in ensuring equitable diagnostic performance. Our contributions can be summarized as follows:
@@
  • A model- and task-agnostic evaluation framework with a compact yet comprehensive representation, both qualitatively and quantitatively, of multiple utility-fairness trade-offs resulting from the deployment of ML systems, facilitating system performance analysis and comparison.
@@
  • A detailed analysis of the proposed framework and rationale through simulations of typical ML systems on synthetically generated data.
@@
  • An empirical study based on three real-world medical imaging datasets demonstrating the effectiveness of the proposed framework.
  • An open-source implementation of the framework allowing reproduction of results established in this article, and further reuse [^1] https://pypi.org/project/fairical.
While the framework is demonstrated in the context of medical imaging, its formulation is model-, metric-, and domain-agnostic, and can be readily applied to other high-stakes ML systems (e.g., decision-support in finance, hiring, criminal justice, or homeland security applications for biometrics) where multiple fairness and utility objectives may be in conflict. By providing a unified view of utility-fairness trade-off, the framework establishes a generalizable flow for cross-domain benchmarking and transparent model selection across diverse application areas.
@@
The paper is organized as follows: Section 2 reviews previous works tackling the evaluation of utility-fairness trade-off systems. Section 3 thoroughly describes the proposed evaluation framework, with a set of use cases and MOO principles behind it. The applicability of the framework in real-world scenarios is shown in Section 4, with the analysis of ML systems for three medical imaging tasks. Finally, a discussion and conclusion with the key points and limitations from this study are presented in Section 5 and Section 6.
2 Background and Related Work
@@
Fairness in machine learning can be categorized based on criteria, sources of bias, perspectives, methodologies, and trade-offs. Fairness criteria include demographic parity, equality of opportunity, equalized odds, and predictive parity (Hardt et al., 2016; Agarwal et al., 2018), each focusing on equitable outcomes or error rates across groups. Sources of bias can stem from data (e.g., under-representation) (Garin et al., 2023), algorithms (e.g., prioritizing accuracy/utility over fairness) (Buolamwini and Gebru, 2018), or human involvement (e.g., subjective labeling) (Zhang et al., 2024). Perspectives of fairness include individual fairness (similar individuals receive similar predictions) (Dwork et al., 2012; Petersen et al., 2021), group fairness (equitable treatment across groups) (Diana et al., 2021; Chan et al., 2024), and subgroup fairness (addressing intersectional identities) (Kuratomi et al., 2025). Methodologies to enforce fairness involve pre-processing data (e.g., balancing representation) (Jang and Wang, 2023; Liu et al., 2021; Lahoti et al., 2020), in-processing adjustments (e.g., modifying loss functions) (Jovanović et al., 2023; Roy and Boddeti, 2019), and post-processing predictions (e.g., calibration techniques) (Hardt et al., 2016; Kim et al., 2020; Jang et al., 2022). Achieving fairness often involves trade-offs, such as balancing it with utility (Liu and Vicente, 2022; Wang et al., 2021) and interpretability (Jo et al., 2023).
Fairness-aware evaluation has an increasing attention in medical imaging tasks, where model predictions can directly affect downstream diagnostic decisions. Recent studies have shown that performance can vary across demographic groups due to acquisition protocols, device related differences, or data imbalance, motivating the inclusion of fairness metrics alongside performance metrics. For instance, in glaucoma detection, ML systems have been evaluated not only for accuracy but also for their fairness performance using equity scaling measurements (Luo et al., 2024) or with respect to demographic features such as gender (Akter et al., 2022). Similarly, multi-objective formalizations have been explored to jointly optimize imaging quality, diagnostic performance and fairness (Lu et al., 2020). However, these works typically address a single fairness notion and do not provide a general framework capable of comparing multiple ML systems under several concurrent fairness constraints, which is the gap our framework aims to bridge.
While several tools and frameworks exist for fairness assessment, their scope differs significantly from the evaluation problem addressed in this work. For instance, Fairlearn and its dashboard (Weerts et al., 2023) and FACET (Gustafson et al., 2023) provide analytics for understanding model behavior across demographic groups, but they evaluate individual model performance rather than assessing the aggregated structure of a utility-fairness trade-off system achieved with our framework. This is an important aspect as the assessment of individual models drawn from a hypothesis class $\mathcal{H}$ may be unstable, whereas evaluating the joint behavior of an entire set of models provides a more stable, comprehensive and complete representation of the fairness performance of the ML system in consideration. Little (2023) proposes a metric that summarizes the utility-fairness trade-off along a single fairness criterion; however, it does not generalize to multi-objective scenarios (multiple utility/fairness). Liu and Vicente (2022) model utility-fairness interactions from a multi-objective perspective, but their formalization focuses on optimization rather than evaluation, and does not provide model-agnostic tools for analyzing the geometry or quality of the trade-off optimality. To the best of our knowledge, no existing framework offers a unified, multi-objective evaluation protocol capable of assessing multiple utility and fairness criteria jointly, and this aspect motivates the contribution of our proposed framework.
@@
Whereas ML systems are typically developed (and evaluated) using a single utility criterion, they are often deployed in scenarios where multiple objectives must be respected. A modern example of this condition relates to the deployment of ML systems under one or multiple demographic fairness constraints (Liu and Vicente, 2022; Zhang et al., 2021; Padh et al., 2021). In this context, we argue that evaluation techniques cross-pollinated from multi-objective optimization (MOO) offer a rich set of primitives allowing for a comprehensive performance characterization under multiple criteria that can streamline system evaluation in this realm.
@@
The principal aim of MOO is to find solutions that lie on, or are proximate to, the set of the optimal performance points called the Pareto Front (PF), resulting in a spectrum of ideal trade-offs among the various objectives. This methodology equips decision-makers with the means to select the most favorable compromise amidst conflicting goals, fostering more informed and balanced decision-making (Wu and Azarm, 2001). The trade-off selection procedure in MOO is therefore critical and affects the quality of service for the deployed system, especially in the case of conflicting objectives. Assessing the quality of these trade-off systems is comparative and encompasses criteria such as proximity to the Pareto optimal set (convergence), the distribution/spread of the points in the objective space (diversity), and the cardinality of solutions (capacity) (Zitzler et al., 2003). These criteria are evaluated by MOO specific performance indicators that have been studied in previous works (refer to Section 3.2.1 for details) (Tan et al., 2002; Wu and Azarm, 2001; Van Veldhuizen and Lamont, 2000; Coello Coello and Reyes Sierra, 2004).
@@
Even though the modeling of trade-off for demographic fairness-accuracy PF is well-known (Wei and Niethammer, 2022; Zietlow et al., 2022), performance indicators for the quality of the PF have rarely been exploited in the context of fairness. Yang et al. (2023) developed a bias mitigation framework that incorporated the Area Under the Curve (AUC) metric, while considering both inter- and intra-group AUC simultaneously. However, the bias mitigation framework does not provide an evaluation protocol for the utility-fairness trade-off; instead, it leverages the AUC to address fairness performance. Little (2023) proposed a scalar measure of the area under the curve from the trade-off between fairness and accuracy. The generated curve outlines the empirical Pareto frontier consisting of the highest attained accuracy within a collection of fitted models at every level of fairness. Although Little (2023) focuses on similar issues as in this study, it does not address the challenge of comparing multiple ML systems in high dimensions. Additionally, the analysis of the PF is superficial, ignoring important performance indicators for diversity and capacity, providing an incomplete evaluation of compared ML strategies. To tackle the aforementioned issues, a more flexible evaluation framework is needed to accommodate different fairness criteria and utility metrics, facilitating a straightforward comparison and analysis of results from different algorithms. The method should be model-agnostic, allowing for real-world comparisons among trade-off systems that may have been optimized using different objectives. Furthermore, since the utility goals of the model across multiple objectives often diverge from fairness goals, performance indicators of the optimal PF solutions can provide a deeper understanding of the trade-offs across these objectives (Wang et al., 2021). The proposed evaluation framework bridges the gap between these issues and their solutions, providing a comprehensive guideline on how this can be achieved in the following sections.
While prior works typically quantify fairness outcomes into a scalar metric or report the best performing ML system, such simplifications may create ambiguity on the relationship between different demographic attributes. The evaluation perspective considered in this study bridges this gap by encouraging analysis across the full spectrum of utility-fairness trade-offs.
3 Methodology
@@
In MOO, each individual objective is considered a distinct characteristic that needs to be optimized to its fullest potential. The trajectory of the optimization is determined by the objective functions in a cooperative way. Conflicting objectives increase the complexity of optimization forcing cooperation, thus making it harder to achieve optimal solutions. This trade-off is measured at evaluation time by using multiple metrics directly related to each of the objectives. Likewise, utility and fairness can result in conflicting objectives challenging the optimization of both through the same ML strategies. A good level of utility is typically achieved by sacrificing fairness or through less biased models that might reduce utility.
@@
Figure 1: Black-box system evaluation (Scenario 1).
3.1 Use Cases